The diagonal of a square is 7 radical 2. What is the length of each side?
Answer by checkley77(12844)

X^2+X^2=(7SQRT2)^2
2X^2=49*2
2X^2=98
X^2=98/2
X^2=49
X=SQRT49
X=7 ANS. FOR THE LENGTH OF THE SIDES OF THE SQUARE.